# Nonlinear Causal Discovery through a Sequential Edge Orientation Approach

### v1

@@ 2.3 Additive Noise Models
There are recent developments on learning causal DAGs from observational data, assuming the additive noise model (ANM) (Hoyer et al., 2008). Under the ANM, each variable $X_{i}$ is a function of its parents nodes $\text{PA}_{i}$ in DAG $\mathcal{G}_{0}$ plus an independent additive noise $\varepsilon_{i}$, i.e.
@@
where $f_{i}$ is an arbitrary function for each $i$ and the noise variables are jointly independent. Moreover, the parents and the noise term are independent of each other, i.e. $\text{PA}_{i}\perp\!\!\!\!\perp\varepsilon_{i}$. A restricted additive noise model is an ANM with restrictions on the functions $f_{i}$, conditional distributions of $X_{i}$, and noise variables (Peters et al., 2014). In particular, the functions must be three-times continuously differentiable. See Definition 3 in the Appendix for the full definition. Throughout this work, we assume causal minimality for the ANM, which is satisfied as long as the function $f_{i}$, for all $i$, is not constant in any of its arguments (Peters et al., 2014). Under the causal minimality assumption, a key result is that the true DAG $\mathcal{G}_{0}$ can be identified from the joint distribution $p(X_{1},\dots,X_{p})$ when the SEM for $\{X_{i}\}_{i=1}^{p}$ satisfies a restricted additive noise model.
We also consider the causal additive model (CAM), a special case of the ANM where both the functions and the error term are additive in the SEM (Bühlmann et al., 2014). The model is defined as
@@
where functions $f_{i,j}$ are three times differentiable and the error terms $\varepsilon_{1},\dots,\varepsilon_{p}$ are independent and $\varepsilon_{i}\sim N(0,\sigma_{i}^{2})$ with $\sigma_{i}^{2}>0$. Bühlmann et al. (2014) have demonstrated that the true DAG $\mathcal{G}_{0}$ can be identified from the joint distribution $p(X_{1},\dots,X_{p})$ when the SEM for $\{X_{i}\}_{i=1}^{p}$ satisfies a CAM. Moreover, the source nodes are allowed to have a non-Gaussian density. As the true SEM under an ANM is difficult to recover in a practical setting, assuming a CAM assists in better recovering the underlying causal relations, by utilizing regression models such as the generalized additive model.
We further assume that the nonlinear functions $f_{i,j}$ belong to a function class $\mathcal{F}$ of smooth functions that is closed with respect to the $\mathcal{L}_{2}$ norm of $P(X_{i}),i=1,\ldots,p$. Let $b_{j},j=1,\dots,m_{n}$ be a set of basis functions, with $m_{n}\rightarrow\infty$ sufficiently slowly as sample size $n\rightarrow\infty$. We then define the general space of additive functions $\mathcal{F}_{n}$ and the space of additive functions $\mathcal{F}_{n}^{\oplus k}$ with respect to finite dimension $k$ as
@@ 3 Algorithm Overview
Classical causal discovery methods learn an equivalence class of the true DAG, represented by a CPDAG. Since nonlinear DAG models are identifiable (Peters et al., 2014), we aim to further infer the causal directions of undirected edges in a CPDAG or, more generally, a PDAG for the finite sample case. To this end, we formulate a novel criterion, named the pairwise additive noise model (PANM), to determine which undirected edges in a PDAG can be oriented at the current stage in a sequential manner. Once an edge fulfilling this criterion is oriented, we further utilize additional graphical rules to orient more undirected edges. We show that this sequential procedure is able to orient all undirected edges in a CPDAG into the true causal DAG, thus accomplishing the goal of identifying all causal relations among the variables. In this section, we introduce the pairwise additive noise model, the sequential edge orientation procedure, and the finite sample version of our full algorithm.
@@ 3.1 Pairwise Additive Noise Model
The pairwise additive noise model, a set of SEMs defined with respect to two nodes, encapsulates the conditions sufficient for edge orientation. For an undirected edge $X-Y$ in a PDAG $\mathcal{G}$, given their parent sets $\text{pa}_{\mathcal{G}}(X),\text{pa}_{\mathcal{G}}(Y)$, its causal direction in true DAG $\mathcal{G}_{0}$ can be identified when $(X,Y)$ follows a PANM. In certain cases, a pair of nodes may follow a PANM even if their parent sets in $\mathcal{G}_{0}$ are not fully identified, i.e. $\text{pa}_{\mathcal{G}}(v)\subset\text{pa}_{\mathcal{G}_{0}}(v)$. We demonstrate how this criterion lays the foundation for our edge orientation procedure.
@@
Let $X,Y$ be two random variables and $Z_{1},Z_{2}$ be two sets of random variables. We say that $[X,Y\mid Z_{1},Z_{2}]$ follows a pairwise additive noise model if either (i) or (ii) holds:
@@
Within the context of causal learning, the structural equation models in conditions (i) and (ii) correspond to the additive noise models under $X\rightarrow Y$ and $Y\rightarrow X$, respectively, where $Z_{1}=\text{pa}_{\mathcal{G}}(X)$ and $Z_{2}=\text{pa}_{\mathcal{G}}(Y)$ in a PDAG $\mathcal{G}$. In this regard, we may simply say that the undirected edge $X-Y$ or the pair of nodes $(X,Y)$ in the PDAG $\mathcal{G}$ satisfies the PANM. To determine whether $(X,Y)$ follows a PANM, we can test if the independence relations hold. More specifically, only the PANM constructed under the correct orientation of $X-Y$ yields the independence relations between the noise term and the parent variables. The precise statement is summarized into the following identifiability result:
@@
Assume $\{X_{i}\}_{i=1}^{p}$ follows a restricted ANM with respect to a DAG $\mathcal{G}_{0}$. Suppose two nodes $X,Y$ are connected by an undirected edge in a PDAG $\mathcal{G}$ which has a consistent extension to $\mathcal{G}_{0}$. If $[X,Y|\text{pa}_{\mathcal{G}}(X),\text{pa}_{\mathcal{G}}(Y)]$ follows the pairwise ANM, then the causal direction between $X$ and $Y$ is identifiable.
@@
The connection between the PANM and the identifiability of the causal direction is more concretely illustrated in Figure 1, showing cases when the true causal direction can and cannot be recovered. Depicted in example (a), edge $X-Y$ can be correctly oriented because the causal relations depicted in the DAG (top) yield the correctly specified pairwise ANM and the true parent sets $\text{pa}_{\mathcal{G}}(X)=\text{pa}_{\mathcal{G}}(Y)=\{A,B\}$, except that between $X$ and $Y$, are identified in the PDAG (bottom). The independence relations hold for SEMs built according to $X\rightarrow Y$. Panel (b) shows an example in which an undirected edge satisfies the PANM even though the parent sets are not fully identified. The SEMs corresponding to $X\rightarrow Y$ are $Y=f_{Y}(X,B)+\epsilon_{Y}$, which is the true SEM, and $X=\widetilde{\epsilon}_{X}$, where $\widetilde{\epsilon}_{X}=g(A,B)+\epsilon_{X}$. While the parents $A,B$ for $X$ are not detected in the PDAG as $\text{pa}_{\mathcal{G}}(X)=\varnothing$, they are merged with the error term $\epsilon_{X}$ to form the new error $\widetilde{\epsilon}_{X}$. Thus, $X-Y$ satisfies the PANM and its orientation is identifiable.
$A$$B$$X$$Y$$Z$
$A$$B$$X$$Y$
$A$$B$$Z$$X$$Y$
$X$$A$$Y$
$A$$B$$X$$Y$$Z$(a)
$A$$B$$X$$Y$(b)
$A$$B$$Z$$X$$Y$(c)
$X$$A$$Y$(d)
@@
Figure 1: Examples to illustrate the PANM. The top row shows the true DAG, while the bottom row features a PDAG extendable to the true DAG with the evaluated edge $X-Y$. (a) $(X,Y)$ satisfy the PANM because both parent sets are fully identified. (b) $(X,Y)$ satisfy the PANM, despite $A,B$ missing from $\text{pa}_{\mathcal{G}}(X)$, since $A,B$ can be merged with $\epsilon_{X}$ to represent $X=\widetilde{\epsilon}_{X}=g(A,B)+\epsilon_{X}$. (c) $[X,Y\mid A,B]$ does not form a PANM, as common parent $Z$ is not detected and becomes a latent confounder in the model. (d) $(X,Y)$ do not satisfy the PANM since node $Y$ is missing parent $A$, which does not guarantee $\varepsilon_{Y}\perp\!\!\!\!\perp X$.
@@
However in example (c), the common parent node $Z$ is not identified in the PDAG. The SEMs constructed over $[X,Y\mid A,B]$ would not satisfy either set of independence relations in Definition (ii) due to the presence of a hidden confounder $Z$, which would result in incorrect or inconclusive inference on the causal direction. In some cases, the independence relations entailed by the PANM may not hold if a parent node, say of $Y$, on a directed path from $X$ to $Y$ is not identified. Consider the undirected edge $X-Y$ in example (d), where $\text{pa}_{\mathcal{G}}(X)=\text{pa}_{\mathcal{G}}(Y)=\varnothing$ in the PDAG. Node $A$ is a parent of $Y$ in the DAG $\mathcal{G}_{0}$ on the directed path $X\to A\to Y$, but it is not identified as a parent of $Y$ in the PDAG. We now examine whether the model $[X,Y]$ satisfies PANM under $X\rightarrow Y$. Node $X$ has no parents in the true DAG, hence its SEM $X=\varepsilon_{X}$ matches the true form. Yet, node $Y$ can only be expressed as $Y=f_{Y}(X,A)+\varepsilon_{Y}=f_{Y}(X,g(X)+\varepsilon_{A})+\varepsilon_{Y}$, where $A$ is substituted as $A=g(X)+\varepsilon_{A}$ and marginalized out. This shows that the SEM for $Y$ is not an additive noise model. Suppose $\widehat{Y}$ is the best approximation of $Y$ by only functions of $X$. The residual $Y-\widehat{Y}$ then in general depends on $X$, so the independence relation $(Y-\widehat{Y})\perp\!\!\!\!\perp X$ does not hold.
@@ 3.2 Key Idea: Sequential Orientation of Edges
We present the high-level, population version of the sequential edge orientation procedure in Algorithm 1 to demonstrate this core idea. Given a PDAG, the algorithm identifies an edge $(i,j)$ which satisfies the PANM on Line 1 and then orients the edge into its true causal direction which can be identified (Lemma 1). Subsequently, the algorithm leverages information read from $\mathcal{G}$ to further identify common children of $i$ and $j$ on Line 1. We denote the set of neighbors and children of $i$ as $\text{nc}_{\mathcal{G}}(i):=\text{ne}_{\mathcal{G}}(i)\cup\text{ch}_{\mathcal{G}}(i)$. For each node $k\in\text{nc}_{\mathcal{G}}(i)\cap\text{nc}_{\mathcal{G}}(j)$, we orient the edges $i\rightarrow k$ and $j\rightarrow k$. Figure 2 features three scenarios in which $\text{nc}_{\mathcal{G}}(i)\cap\text{nc}_{\mathcal{G}}(j)\neq\varnothing$ occurs and Line 1 is applied: (1) $k\in\text{ne}_{\mathcal{G}}(i)\cap\text{ne}_{\mathcal{G}}(j)$, (2) $k\in\text{ne}_{\mathcal{G}}(i)\cap\text{ch}_{\mathcal{G}}(j)$, and (3) $k\in\text{ch}_{\mathcal{G}}(i)\cap\text{ne}_{\mathcal{G}}(j)$. Case 2 and case 3 respectively correspond to when $j\rightarrow k$ and $i\rightarrow k$ have been oriented by prior actions before evaluating $i-j$. For all three cases, we orient $k$ as a common child of $i$ and $j$ as shown in the bottom panel of Figure 2. This is because of the following reasoning: If $k$ is a common parent node of $i,j$, then $[i,j\mid\text{pa}_{\mathcal{G}}(i),\text{pa}_{\mathcal{G}}(j)]$ would not satisfy the independence relations entailed by the PANM as $k$ would be a hidden confounder, as discussed in Figure 1c. If the true orientation were $i\to k\to j$ or $j\to k\to i$, this would be the case of Figure 1d and again would violate the PANM assumptions. Therefore, $k$ must be a child node of both $i$ and $j$. On the following line, the procedure applies rule 1 of Meek’s orientation rules, where the configuration $a\rightarrow b-c$ is oriented as $a\rightarrow b\rightarrow c$ given there is no edge between $a$ and $c$, to identify descendant nodes of $a$. Finally, if no undirected edges satisfy the condition on Line 1, we apply the Meek’s rules to maximally orient the PDAG on Line 1.
@@
3       Search for $(i,j)\in U$ such that $[i,j\mid\text{pa}_{\mathcal{G}}(i),\text{pa}_{\mathcal{G}}(j)]$ satisfies the PANM;

### v2

@@ 2.3 Additive Noise Models
There are recent developments on learning causal DAGs from observational data, assuming the additive noise model (ANM) (hoyer2008nonlinear). Under the ANM, each variable $X_{i}$ is a function of its parent nodes $\text{PA}_{i}$ in DAG $\mathcal{G}_{0}$ plus an independent additive noise $\varepsilon_{i}$, i.e.
@@
where $f_{i}$ is an arbitrary function for each $i$ and the noise variables are jointly independent. Moreover, the parents and the noise term are independent of each other, i.e. $\text{PA}_{i}\perp\!\!\!\!\perp\varepsilon_{i}$. A restricted additive noise model is an ANM with restrictions on the functions $f_{i}$, conditional distributions of $X_{i}$, and noise variables (resit_anm). In particular, the functions must be three-times continuously differentiable. See Definition 3 in the Appendix for the full definition. Throughout this work, we assume causal minimality for the ANM, which is satisfied as long as the function $f_{i}$, for all $i$, is not constant in any of its arguments (resit_anm). Under the causal minimality assumption, a key result is that the true DAG $\mathcal{G}_{0}$ can be identified from the joint distribution $p(X_{1},\dots,X_{p})$ when the SEM for $\{X_{i}\}_{i=1}^{p}$ satisfies a restricted additive noise model.
We also consider the causal additive model (CAM), a special case of the ANM where both the function $f_{i}$ and the error term are additive in the SEM (cam). The model is defined as
@@
where functions $f_{i,j}$ are three times differentiable and the error terms $\varepsilon_{1},\dots,\varepsilon_{p}$ are independent and $\varepsilon_{i}\sim N(0,\sigma_{i}^{2})$ with $\sigma_{i}^{2}>0$. cam have demonstrated that the true DAG $\mathcal{G}_{0}$ can be identified from the joint distribution $p(X_{1},\dots,X_{p})$ when the SEM for $\{X_{i}\}_{i=1}^{p}$ satisfies a CAM. Moreover, the source nodes are allowed to have a non-Gaussian density. As the true SEM under a general ANM is difficult to recover in a practical setting, assuming a CAM assists in better recovering the underlying causal relations, by utilizing regression models such as the generalized additive model.
We further assume that the nonlinear functions $f_{i,j}$ belong to a function class $\mathcal{F}$ of smooth functions that is closed with respect to the $\mathcal{L}_{2}$ norm of $P(X_{i}),i=1,\ldots,p$. Let $b_{j},j=1,\dots,m_{n}$ be a set of basis functions, with $m_{n}\rightarrow\infty$ sufficiently slowly as sample size $n\rightarrow\infty$. We then define the general function space $\mathcal{F}_{n}$ and the space of additive functions $\mathcal{F}_{n}^{\oplus k}$ with respect to finite dimension $k$ as
@@ 3 Algorithm Overview
Classical causal discovery methods learn an equivalence class of the true DAG, represented by a CPDAG. Since nonlinear DAG models are identifiable (resit_anm), we aim to further infer the causal directions of undirected edges in a CPDAG or, more generally, a PDAG for the finite sample case. To this end, we formulate a novel criterion, named the pairwise additive noise model (PANM), to determine which undirected edges in a PDAG can be oriented at the current stage in a sequential manner. Once an edge fulfilling this criterion is oriented, we further utilize additional graphical rules to orient more undirected edges. We show that this sequential procedure is able to orient all undirected edges in a CPDAG into the true causal DAG, thus accomplishing the goal of identifying all causal relations among the variables. In this section, we introduce the pairwise additive noise model, the sequential edge orientation procedure, and the finite sample version of our full algorithm.
@@ 3.1 Pairwise Additive Noise Model
The pairwise additive noise model, a set of SEMs defined with respect to two nodes, encapsulates the conditions sufficient for edge orientation. For an undirected edge XYX-Y in a PDAG $\mathcal{G}$, given their parent sets $\text{pa}_{\mathcal{G}}(X)$ and $\text{pa}_{\mathcal{G}}(Y)$, its causal direction in true DAG $\mathcal{G}_{0}$ can be identified when $(X,Y)$ follows a PANM. In certain cases, a pair of nodes may follow a PANM even if their parent sets in $\mathcal{G}_{0}$ are not fully identified, i.e. $\text{pa}_{\mathcal{G}}(v)\subset\text{pa}_{\mathcal{G}_{0}}(v)$. We demonstrate how this criterion lays the foundation for our edge orientation procedure.
@@
Let $X,Y$ be two random variables and $Z_{1},Z_{2}$ be two sets of random variables. We say that $X,Y\mid Z_{1},Z_{2}]$ follows a pairwise additive noise model if either (i) or (ii) holds:
@@
Furthermore, we assume both SEMs above satisfy Condition 1 in Appendix A for any value $(z_{1},z_{2})$ in the domain of $(Z_{1},Z_{2})$.
Within the context of causal learning, the structural equation models in conditions (i) and (ii) correspond to the additive noise models under $X\rightarrow Y$ and $Y\rightarrow X$, respectively, where $Z_{1}=\text{pa}_{\mathcal{G}}(X)$ and $Z_{2}=\text{pa}_{\mathcal{G}}(Y)$ in a PDAG $\mathcal{G}$. In this regard, we may simply say that the undirected edge XYX-Y or the pair of nodes $(X,Y)$ in the PDAG $\mathcal{G}$ satisfies the PANM. To determine whether $(X,Y)$ follows a PANM, we can test if the independence relations hold. More specifically, only the PANM constructed under the correct orientation of XYX-Y yields the independence relations between the noise term and the parent variables. The precise statement is summarized into the following lemma, which is an immediate consequence of the identifiability of bivariate ANMs (hoyer2008nonlinear).
@@
Assume $\{X_{i}\}_{i=1}^{p}$ follows a restricted ANM with respect to a DAG $\mathcal{G}_{0}$. Suppose two nodes $X,Y$ are connected by an undirected edge in a PDAG $\mathcal{G}$ which has a consistent extension to $\mathcal{G}_{0}$. If $X,Y|\text{pa}_{\mathcal{G}}(X),\text{pa}_{\mathcal{G}}(Y)]$ follows the PANM, then the causal direction between $X$ and $Y$ is identifiable.
@@
The connection between the PANM and the identifiability of the causal direction is more concretely illustrated in Figure 1, showing cases when the true causal direction can and cannot be recovered. Depicted in example (a), edge XYX-Y can be correctly oriented because the causal relations depicted in the DAG (top) yield the correctly specified pairwise ANM and the true parent sets $\text{pa}_{\mathcal{G}}(X)=\text{pa}_{\mathcal{G}}(Y)=\{A,B\}$, except that between $X$ and $Y$, are identified in the PDAG (bottom). The independence relations hold for SEMs built according to $X\rightarrow Y$. Panel (b) shows an example in which an undirected edge satisfies the PANM even though the parent sets are not fully identified. The SEMs corresponding to $X\rightarrow Y$ are $Y=f_{Y}(X,B)+\varepsilon_{Y}$, which is the true SEM, and $X=\widetilde{\varepsilon}_{X}$, where $\widetilde{\varepsilon}_{X}=g(A,B)+\varepsilon_{X}$. While the parents $A,B$ for $X$ are not detected in the PDAG as $\text{pa}_{\mathcal{G}}(X)=\varnothing$, they are merged with the error term $\varepsilon_{X}$ to form the new error $\widetilde{\varepsilon}_{X}$. Thus, XYX-Y satisfies the PANM and its orientation is identifiable.
AABBXXYYZZ
AABBXXYY
AABBZZXXYY
XXAAYY
AABBXXYYZZ(a)
AABBXXYY(b)
AABBZZXXYY(c)
XXAAYY(d)
@@
Figure 1: Examples to illustrate the PANM. The top row shows the true DAG, while the bottom row features a PDAG extendable to the true DAG with the evaluated edge XYX-Y. (a) $X,Y\mid A,B]$ satisfies the PANM because both parent sets are fully identified. (b) $X,Y\mid\varnothing,B]$ satisfies the PANM, despite $A,B$ missing from $\text{pa}_{\mathcal{G}}(X)=\varnothing$, since we can write $X=\widetilde{\varepsilon}_{X}=g(A,B)+\varepsilon_{X}$. (c) $X,Y\mid A,B]$ does not form a PANM, as common parent $Z$ is not detected and becomes a latent confounder in the model. (d) $[X,Y]$ does not satisfy the PANM since node $Y$ is missing parent AA, which does not guarantee $\varepsilon_{Y}\perp\!\!\!\!\perp X$.
@@
In example (c), however, the common parent node $Z$ is not identified in the PDAG. The SEMs constructed over $X,Y\mid A,B]$ would not satisfy either set of independence relations in Definition (ii) due to the presence of a hidden confounder $Z$, which would result in incorrect or inconclusive inference on the causal direction. In some cases, the independence relations entailed by the PANM may not hold if a parent node, say of $Y$, on a directed path from $X$ to $Y$ is not identified. Consider the undirected edge XYX-Y in example (d), where $\text{pa}_{\mathcal{G}}(X)=\text{pa}_{\mathcal{G}}(Y)=\varnothing$ in the PDAG. Node $A$ is a parent of $Y$ in the DAG $\mathcal{G}_{0}$ on the directed path $X\to A\to Y$, but it is not identified as a parent of $Y$ in the PDAG. We now examine whether the model $[X,Y]$ satisfies PANM under $X\rightarrow Y$. Node $X$ has no parents in the true DAG, hence its SEM $X=\varepsilon_{X}$ matches the true form. Yet, node $Y$ can only be expressed as $Y=f_{Y}(X,A)+\varepsilon_{Y}=f_{Y}(X,g(X)+\varepsilon_{A})+\varepsilon_{Y}$, where $A$ is substituted as $A=g(X)+\varepsilon_{A}$ and marginalized out. This shows that the SEM for $Y$ is not an additive noise model. Suppose $\widehat{Y}$ is the best approximation of $Y$ by only functions of $X$. Then, the residual $Y-\widehat{Y}$ in general depends on $X$, so the independence relation $(Y-\widehat{Y})\perp\!\!\!\!\perp X$ does not hold.
@@ 3.2 Key Idea: Sequential Orientation of Edges
We present the high-level, population version of the sequential edge orientation procedure in Algorithm 1 to demonstrate this core idea. Given a PDAG, the algorithm identifies an edge $(i,j)$ which satisfies the PANM on Line 1 and then orients the edge into its true causal direction which can be identified (Lemma 1). Subsequently, the algorithm leverages information read from $\mathcal{G}$ to further identify common children of $i$ and $j$ on Line 1. We denote the set of neighbors and children of $i$ as $\text{nc}_{\mathcal{G}}(i):=\text{ne}_{\mathcal{G}}(i)\cup\text{ch}_{\mathcal{G}}(i)$. For each node $k\in\text{nc}_{\mathcal{G}}(i)\cap\text{nc}_{\mathcal{G}}(j)$, we orient the edges $i\rightarrow k$ and $j\rightarrow k$. Figure 2 features three scenarios in which $\text{nc}_{\mathcal{G}}(i)\cap\text{nc}_{\mathcal{G}}(j)\neq\varnothing$ occurs and Line 1 is applied: (1) $k\in\text{ne}_{\mathcal{G}}(i)\cap\text{ne}_{\mathcal{G}}(j)$, (2) $k\in\text{ne}_{\mathcal{G}}(i)\cap\text{ch}_{\mathcal{G}}(j)$, and (3) $k\in\text{ch}_{\mathcal{G}}(i)\cap\text{ne}_{\mathcal{G}}(j)$. Case 2 and case 3 respectively correspond to when $j\rightarrow k$ and $i\rightarrow k$ have been oriented by prior actions before evaluating iji-j. For all three cases, we orient $k$ as a common child of $i$ and $j$ as shown in the bottom panel of Figure 2. This is because of the following reasoning: If $k$ were a common parent node of $i,j$, then $i,j\mid\text{pa}_{\mathcal{G}}(i),\text{pa}_{\mathcal{G}}(j)]$ would not satisfy the independence relations entailed by the PANM since $k$ would be a hidden confounder, as discussed in Figure 1c. If the true orientation were $i\to k\to j$ or $j\to k\to i$, this would be the case of Figure 1d and again would violate the PANM assumptions. Therefore, $k$ must be a child node of both $i$ and $j$. On the following line, the procedure applies rule 1 of Meek’s orientation rules, where the configuration $a\rightarrow b-c$ is oriented as $a\rightarrow b\rightarrow c$ given there is no edge between $a$ and $c$, to identify descendant nodes of $a$. Finally, if no undirected edges satisfy the condition on Line 1, we apply the Meek’s rules to maximally orient the PDAG on Line 1.
@@
3   Search for $(i,j)\in U$ such that $i,j\mid\text{pa}_{\mathcal{G}}(i),\text{pa}_{\mathcal{G}}(j)]$ satisfies the PANM;